# DDPG-E2E: A Novel Policy Gradient Approach for End-to-End Communication Systems

### v1

@@ Abstract
The End-to-end (E2E) learning-based approach has great potential to reshape the existing communication systems by replacing the transceivers with deep neural networks. To this end, the E2E learning approach needs to assume the availability of prior channel information to mathematically formulate a differentiable channel layer for the backpropagation (BP) of the error gradients, thereby jointly optimizing the transmitter and the receiver. However, accurate and instantaneous channel state information is hardly obtained in practical wireless communication scenarios. Moreover, the existing E2E learning-based solutions exhibit limited performance in data transmissions with large block lengths. In this article, these practical issues are addressed by our proposed deep deterministic policy gradient-based E2E communication system. In particular, the proposed solution utilizes a reward feedback mechanism to train both the transmitter and the receiver, which alleviates the information loss of error gradients during BP. In addition, a convolutional neural network (CNN)-based architecture is developed to mitigate the curse of dimensionality problem when transmitting messages with large block lengths. Extensive simulations then demonstrate that our proposed solution can not only jointly train the transmitter and the receiver simultaneously without requiring the prior channel knowledge but also can obtain significant performance improvement on block error rate compared to state-of-the-art solutions.
@@ I Introduction
In the conventional communication system, the data transmission entails multiple signal processing blocks where each module is separately designed and optimized for a specific task, as illustrated in Fig. 1(a). Such a multi-block architecture has achieved reasonable performance, but at the cost of increasing the design complexity. The reason is that it requires careful fine-tuning and optimizations for every individual module with different objectives. Consequently, it is difficult, if not impossible, to jointly optimize the transmitter and the receiver [2]. In addition, these mathematical formulated modules cannot always coherently and accurately capture the actual communication conditions, thereby compromising the data transmission performance [3].
@@
Recently, deep learning (DL) technique has attracted significant attention in the field of communication systems as it provides pure data-driven solutions for handling the imperfection of communication systems. The DL-based approach interprets the end-to-end (E2E) communication system as a deep neural network (DNN)-based auto-encoder over an interference channel [4], where both the transmitter and the receiver are represented as DNNs, as depicted in Fig. 1(b). This E2E-based learning method theoretically assumes the availability of an explicit channel model, which is perceived as an intermediate layer connecting the transmitter and the receiver. The transmitter DNN learns to map the transmitted data into an encoded vector in a smaller dimension and then sends it over the channel. The receiver DNN learns to map the received signals to the estimated data. The E2E system is trained in a supervised manner to minimize its E2E loss which measures the difference between the estimated data and the transmitted data. Thus, this E2E scheme has significantly alleviated the overall design complexity by interpreting the signal processing blocks as DNNs [5]. The E2E paradigm has demonstrated effective performance in simple communication scenarios, e.g., Additive White Gaussian Noise (AWGN) channel [4], while it remains challenging to apply such an E2E system to more complex wireless communication channels, such as frequency-selective channels and Rayleigh block fading channels.
@@
The most critical shortcoming of the E2E-based learning approach is that it requires the prior information of the channel between the transmitter and the receiver. Moreover, it only performs well with differentiable channel models [6]. However, in practical scenarios, the communication channel is often regarded as a black box due to the non-differentiable nature of its transfer function. This characteristic poses a challenge for estimating channel gradients, especially in the context of wireless communication environments. Another critical issue with the E2E learning approach is the transmitter DNN cannot be directly trained via the BP algorithm when the channel transfer function is non-differentiable. During the E2E optimization, the gradient at the receiver can be directly calculated, while the gradient at the transmitter is unavailable due to the non-differentiable channel, which hinders the calculation of the derivative of the loss function at the channel layer. Consequently, it prevents the overall realization of the E2E learning of the communication systems. In addition, the existing E2E learning scheme suffers from the curse of dimensionality. Prior works studying E2E paradigm predominantly focus on the experiments with small block length (e.g., 8 information bits [4]). This occurs because the size of possible input messages grows exponentially with increasing block length when using one-hot encoding. The portion of unseen input message would therefore exponentially increase during the training which leads to significant performance degradation. Research in [7] shows that the E2E paradigm experiences performance degradation, even when only 10% of unseen data is omitted from the training set. Considering the aforementioned challenges, there is a research gap in the E2E learning-based data transmission with large block length over an unknown channel model, while concurrently addressing the problem of local optimization arising from the non-differentiable channel function. Hence, it demands an innovative E2E communication system capable of adaptive training on various channel models without prior knowledge of explicit channel effects. Additionally, the proposed scheme should demonstrate the ability to transmit long sequences of bit-stream without compromising performance.
@@
Given the above, this article introduces a novel deep reinforcement learning (DRL)-based E2E communication system empowered by the deep deterministic policy gradient (DDPG) algorithm, where both the transmitter and the receiver can be jointly trained over an unknown channel. This approach leverages the loss information obtained from the receiver as the reward signal and directly applies it to optimize the transmitter which acts as the agent in the DDPG algorithm. Additionally, a CNN-based architecture is developed to accommodate input and output message with large block length. The proposed scheme exhibits notable improvement in BLER performance compared to existing solutions under complex wireless communication scenarios, particularly when operating with large block lengths (e.g., 256 information bits). Our major contributions are summarized as follows.
@@
  • To address the issue of dimensionality, CNN are employed to extend the block length of the system from 8 bits to 256 bits. This approach allows the transmitter to process high dimensional input data, and the receiver then learns to recover the encoded signals through our proposed CNN framework.
  • Through extensive simulations, we demonstrate that the proposed DDPG-based E2E communication system can achieve better BLER performance and convergence rate compared to state-of-the-art solutions over complex wireless communication channels, e.g., Rayleigh fading and Rician fading channels.
The rest of the paper is organized as follows. The related works will be reviewed in Section II. In Section III, the modeling of E2E communication systems is presented. In Section IV, the proposed DDPG-based E2E communication system is explained in detail. Then, the training process of the E2E system is presented in Section V. The simulation results are presented in Section VI. Finally, Section VII concludes our paper.
Notation: The notations used in this paper are listed in the following. Lowercase letters $\bm{m}$ and $\hat{\bm{m}}$ denote the original message and the estimated message. $\bm{x}$ denotes the encoded signal, and $\bm{y}$ denotes the received signal. ${\mathbb{C}}^{n}$ denotes the complex values of $n$ dimension. $\bm{s}_{t}$, $\bm{a}_{t}$ and $r_{t}$ denote the state, action and reward at time step $t$, respectively. $\mu$ denotes the policy function mapping state $\bm{s}$ to action $\bm{a}$. $Q(\bm{s},\bm{a})$ denotes the Q-value which measures expected cumulative reward that an agent can obtain by taking a particular action $\bm{a}$ in a particular state $\bm{s}$ following a policy function $\mu$.
@@ II Related Works
In this section, existing literature about E2E learning for communication systems to deal with unknown channel models are reviewed.
@@ II-A Deep Learning in End-to-End Communication Systems
In recent years, different DL techniques are applied to E2E communication systems to solve the unknown channel problem. The study is pioneered in [2], a two-step training scheme using transfer learning is developed, where the E2E communication system is trained on an assumed stochastic channel model that is close to the actual channel effects at the first stage. Subsequently, the receiver is then finetuned based on the difference between the assumed stochastic channel model and practical channel model at the second stage. This proposed scheme showcases the possibility of conducting over-the-air data transmission entirely carried out by NNs. However, the transmitter cannot be optimized during the finetuning of the receiver, which hinders the overall learning of E2E communication system. There are some existing literature studying to improve the two-step training approach, which can be classified into two main categories, i.e., channel imitation approaches and receiver-aided approaches.
@@ II-B Channel Imitation Schemes for End-to-End Communication Systems
In order to learn the channel effects in practical channel models and address the challenges of local optimization caused by BP algorithm during the auto-encoder training [8], the channel imitation scheme using generative adversarial network (GAN) is widely studied [9]. In [10], a novel conditional GAN based channel agnostic E2E communication system is developed to adaptively learn the channel effects of different types of channel models. More specifically, the generator is trained to approximate the conditional data distribution of the practical channel effects, and the discriminator is trained to distinguish the real signals from the generated fake signals. By conditioning the generator and the discriminator on the channel, the system can learn to adapt to a variety of channel conditions, even if the channel characteristics are unknown beforehand. This channel agnostic E2E system empowers the generator to serve as a bridge connecting the transmitter and the receiver. This enables the BP of error gradients from the receiver to the transmitter, ultimately facilitating the overall learning of the transceiver. This conditional GAN based channel imitation scheme is further enhanced in [11], where the system can not only adapt to various channel models by learning their characteristics but can also transmit long sequences of information bits. In particular, this proposed system has a significantly extended block length, from 8 information bits to 128 information bits, without much degradation in BLER performance. However, the conditional GAN based schemes face performance degradation due to the gradient vanishing issue during the BP and the overfitting problem arising from the numerous parameters across the transmitter, receiver, generator and discriminator.
A residual-aided GAN (RA-GAN) based E2E communication system is proposed to solve the above challenges by applying residual neural network (ResNet) [12] to the generator to enable better generation performance for the received signals and adding $\bm{l}_{2}$ regularizer to punish large weights in E2E system [13]. However, this scheme experiments only on a block length of 4 bits, which falls considerably short of the practical scenarios for data transmission. Given the above, although the aforementioned channel imitation schemes have effectively mitigated the local optimization problem caused by BP algorithm, these GAN-based solutions exhibit significant performance degradation with large block length, particularly in more complex wireless channels, such as the Rayleigh fading channel. Further improvements are required to address these challenges and enhance the system’s performance in such scenarios.
@@ II-C Receiver-Aided Schemes for End-to-End Communication Systems
Recently, DRL has demonstrated its great potentials in wireless communications and networking [14], particularly in the field of E2E communication systems design. Apart from the channel imitation solutions, RL techniques have been emerging as a promising tool in receiver-aided approaches, enabling data transmission without the need of explicit channel modelling or CSI [15]. In particular, the simultaneous perturbation stochastic approximation technique is used in [16] to train the transmitter, where the transmitter is optimized by utilizing the loss value obtained from the receiver. In [17], a combination of RL and supervised learning techniques is employed to optimize both the transmitter and the receiver. In [18], the authors extend these works by considering a noisy feedback loop. Particularly, the proposed system with noisy feedback link demonstrates identical performance with the system with perfect feedback link. In summary, the aforementioned receiver-aided solutions circumvent the unknown channel issue and local optimization problem arising from BP by using the loss value obtained from the receiver. However, the biggest issue of the receiver-aided solutions is that the variance of loss value from the receiver will scale with the increasing number of channel uses and the increasing block length of input message, which leads to significant performance degradation and slow convergence. In addition, the training of communication systems over complex wireless communication channels faces challenge due to the relatively long coherence time where only small portions of CSI are learned, which leads to poor sample efficiency and also slow convergence. Given the above disadvantages of the channel imitation approaches and receiver-aided approaches, there is currently no effective solution that can transmit data with large block length (e.g., 256 information bit) over an unknown channel without compromising the BLER performance and the convergence rate.
This article introduces a novel DDPG-based solution that can optimize the E2E systems without prior channel knowledge, and this proposed system can achieve superior BLER and convergence performance in data transmissions with large block length (i.e., 128 bits and 256 bits) under different wireless communication channels, compared to the existing E2E solutions.
@@ III System Model
In this section, we will introduce the basic concepts of E2E communication systems. As shown in Fig. 1(b), an auto-encoder-based E2E communication system consists of a transmitter, a receiver, and a mathematically formulated channel, where the transmitter and the receiver are represented as two independent DNNs. The transmitter DNN aims to encode the original binary message $\bm{m}$ to the encoded signal $\bm{x}\in{\mathbb{C}}^{n}$, where ${\mathbb{C}}^{n}$ denotes that the encoded signal $\bm{x}$ is presented as complex numbers, making $n$ discrete uses of channel. The encoded signal is then sent via the channel to the receiver. The received signal after channel distortion is denoted by $\bm{y}\in{\mathbb{C}}^{n}$. The receiver aims to decode the received signals and recover the original binary message $\bm{m}$. The implementations of the E2E communication system can be illustrated as the sequence of three individual functions, which can be expressed as:
@@
where $f_{\mathrm{E}}$ represents the encoder function mapping the original binary message to the encoded signal, i.e., $\bm{x}=f_{\mathrm{E}}(\bm{m};\bm{\theta}_{\mathrm{E}})$, and $\bm{\theta}_{\mathrm{E}}$ denotes the trainable parameters of the transmitter DNN. $f_{\mathrm{h}}$ represents the channel impairments with channel realization $h$, i.e., $\bm{y}=f_{\mathrm{h}}(\bm{x})$. $f_{\mathrm{D}}$ represents the decoder function that aims to recover the received signals to the estimated message, i.e., $\hat{\bm{m}}=f_{\mathrm{D}}(\bm{y};\bm{\theta}_{\mathrm{D}})$, and $\bm{\theta}_{\mathrm{D}}$ denotes the trainable weights of the receiver DNN. The goal of the E2E communication system is to jointly optimize both the transmitter DNN and the receiver DNN by minimizing the E2E loss function, denoted by $L=\mathcal{L}(\bm{m},\hat{\bm{m}})$. The loss function $\mathcal{L}(\bm{m},\hat{\bm{m}})$ is regarded as the objective function which measures the distance between the original message $\bm{m}$ and the estimated message $\hat{\bm{m}}$ for such E2E framework, and reflects the recovery accuracy of the data transmission.
@@
The E2E learning approach considers the entire system as an auto-encoder and trains the DNNs of the transceivers in a supervised manner. The training process involves calculating the error gradients of each layer by utilizing the BP algorithm. To enable the BP of the error gradients, it is necessary to ensure that the loss function with respect to the trainable weights at each layer of the DNN is differentiable. The computed error gradients subsequently update the weights and bias at each layer, with the goal of minimizing the overall E2E loss. Therefore, the E2E learning approach entails a differentiable channel function beforehand to facilitate the back-propagation of error gradients from the receiver to the transmitter. This enables the joint optimization of the entire E2E communication system. However, while the E2E learning paradigm demonstrates efficiency in simple channel models, its performance becomes notably constrained when confronted by more complex communication scenarios. The current limitations of the E2E learning paradigm can be summarized as follows. First, the primary limitation of the E2E learning method is its reliance on an explicit differentiable channel model, which is often impractical due to real-world communication challenges. Second, the non-differentiable channel model potentially hinders the gradient-based optimization, preventing the overall learning of the system. Third, obtaining real-time CSI is difficult in practice, undermining the validity of assumed channel functions. Furthermore, the existing literature primarily focuses on research with short block lengths, and there is a lack of evidence regarding the performance of solutions for block lengths exceeding 100 bits.
@@
This paper introduces a DDPG-based E2E communication system aimed at addressing the challenges posed by non-differentiable channel models and imperfect CSI, while improving the performance in terms of data transmission with long block lengths. By leveraging the DDPG algorithm, the loss value of the receiver DNN serves as feedback information for updating the transmitter DNN. Hence, the transmitter DNN and the receiver DNN can be implicitly and jointly optimized without requirement of any prior knowledge regarding channel models. The following sections will present comprehensive analysis of our proposed DDPG-based E2E solution.

### v2

@@ Abstract
The End-to-end (E2E) learning-based approach has great potential to reshape the existing communication systems by replacing the transceivers with deep neural networks. To this end, the E2E learning approach needs to assume the availability of prior channel information to mathematically formulate a differentiable channel layer for the back-propagation (BP) of the error gradients, thereby jointly optimizing the transmitter and the receiver. However, accurate and instantaneous channel state information is hardly obtained in practical wireless communication scenarios. Moreover, the existing E2E learning-based solutions exhibit limited performance in data transmissions with large block lengths. In this article, these practical issues are addressed by our proposed deep deterministic policy gradient-based E2E communication system. In particular, the proposed solution utilizes a reward feedback mechanism to train both the transmitter and the receiver, which alleviates the information loss of error gradients during BP. In addition, a convolutional neural network-based architecture is developed to mitigate the curse of dimensionality problem when transmitting messages with large block lengths. Extensive simulations then demonstrate that our proposed solution can not only jointly train the transmitter and the receiver simultaneously without requiring prior channel knowledge but also can obtain significant performance improvement compared to state-of-the-art solutions.
@@ I Introduction
In the conventional communication system, the data transmission entails multiple signal processing blocks where each module is separately designed and optimized for a specific task, as illustrated in Fig. 1(a). Such a multi-block architecture has achieved reasonable performance but at the cost of increasing the design complexity. The reason is that it requires careful fine-tuning and optimizations for every individual module with different objectives. Consequently, it is difficult, if not impossible, to jointly optimize the transmitter and the receiver [2]. In addition, these mathematical formulated modules cannot always coherently and accurately capture the actual communication conditions, thereby compromising the data transmission performance [3].
@@
Recently, deep learning (DL) technique has attracted significant attention in the field of communication systems as it provides pure data-driven solutions for handling the imperfection of communication systems. The DL-based approach interprets the end-to-end (E2E) communication system as a neural network (NN)-based auto-encoder over an interference channel [4], where both the transmitter and the receiver are represented as NNs, as depicted in Fig. 1(b). This E2E-based learning method theoretically assumes the availability of an explicit channel model, which is perceived as an intermediate layer connecting the transmitter and the receiver. The transmitter NN learns to map the transmitted data into an encoded vector in a smaller dimension and then sends it over the channel. The receiver NN learns to map the received signals to the estimated data. The E2E system is trained in a supervised manner to minimize its E2E loss which measures the difference between the estimated data and the transmitted data. Thus, this E2E scheme has significantly alleviated the overall design complexity by interpreting the signal processing blocks as NNs [5]. The E2E paradigm has demonstrated effective performance in simple communication scenarios, e.g., Additive White Gaussian Noise (AWGN) channel [4], while it remains challenging to apply such an E2E system to more complex wireless communication channels, such as frequency-selective channels and Rayleigh block fading channels.
@@
The most critical shortcoming of the E2E-based learning approach is that it requires prior information of the channel between the transmitter and the receiver. Moreover, it only performs well with differentiable channel models [6]. However, in practical scenarios, the communication channel is often regarded as a black box due to the non-differentiable nature of its transfer function. This characteristic poses a challenge for estimating channel gradients, especially in the context of wireless communication environments. Another critical issue with the E2E learning approach is the transmitter NN cannot be directly trained via the BP algorithm when the channel transfer function is non-differentiable. During the E2E optimization, the gradient at the receiver can be directly calculated, while the gradient at the transmitter is unavailable due to the non-differentiable channel, which hinders the calculation of the derivative of the loss function at the channel layer. Consequently, it prevents the overall realization of the E2E learning of the communication systems. In addition, the existing E2E learning scheme suffers from the curse of dimensionality. Prior works studying the E2E paradigm predominantly focus on the experiments with small block lengths (e.g., 8 information bits [4]). This occurs because the size of possible input messages grows exponentially with increasing block length when using one-hot encoding. The portion of unseen input messages would therefore exponentially increase during the training which leads to significant performance degradation. Research in [7] shows that the E2E paradigm experiences performance degradation, even when only 10% of unseen data is omitted from the training set. Considering the aforementioned challenges, there is a research gap in the E2E learning-based data transmission with large block length over an unknown channel model, while concurrently addressing the problem of local optimization arising from the non-differentiable channel function. Hence, it demands an innovative E2E communication system capable of adaptive training on various channel models without prior knowledge of explicit channel effects. Additionally, the proposed scheme should demonstrate the ability to transmit long sequences of bit-stream without compromising performance.
@@
Given the above, this article introduces a novel deep reinforcement learning (DRL)-based E2E communication system empowered by the deep deterministic policy gradient (DDPG) algorithm, where both the transmitter and the receiver can be jointly trained over an unknown channel. This approach leverages the loss information obtained from the receiver as the reward signal and directly applies it to optimize the transmitter which acts as the agent in the DDPG algorithm. Additionally, a CNN-based architecture is developed to accommodate input and output messages with large block lengths. The proposed scheme exhibits notable improvement in BLER performance compared to existing solutions under complex wireless communication scenarios, particularly when operating with large block lengths (e.g., 256 information bits). Our major contributions are summarized as follows.
@@
  • To address the issue of dimensionality, CNN is employed to extend the block length of the system from 8 bits to 256 bits. This approach allows the transmitter to process high dimensional input data, and the receiver then learns to recover the encoded signals through our proposed CNN.
  • Through extensive simulations, we demonstrate that the proposed DDPG-based E2E communication system can achieve better BLER performance and convergence rate compared to state-of-the-art solutions over complex wireless communication channels, e.g., Rician, Rayleigh, and 3GPP channels.
The rest of the paper is organized as follows. The related works will be reviewed in Section II. In Section III, the modeling of E2E communication systems is presented. In Section IV, the proposed DDPG-based E2E system is explained in detail. Then, the training process of the E2E system is presented in Section V. The simulation results are presented in Section VI. Finally, Section VII concludes our paper.
Notation: The notations used in this paper are listed in the following. Lowercase letters $\bm{m}$ and $\hat{\bm{m}}$ denote the original message and the estimated message. $\bm{x}$ denotes the encoded signal, and $\bm{y}$ denotes the received signal. ${\mathbb{C}}^{n}$ denotes the complex values of $n$ dimension. $\bm{s}_{t}$, $\bm{a}_{t}$ and $r_{t}$ denote the state, action and reward at time step $t$, respectively. $\mu$ denotes the policy function mapping state $\bm{s}$ to action $\bm{a}$. $Q(\bm{s},\bm{a})$ denotes the Q-value which measures the expected cumulative reward that an agent can obtain by taking a particular action $\bm{a}$ in a particular state $\bm{s}$ following a policy function $\mu$.
@@ II-A Deep Learning in End-to-End Communication Systems
In recent years, different DL techniques have been applied to E2E communication systems to solve the unknown channel problem. The study is pioneered in [2], a two-step training scheme using transfer learning is developed, where the E2E communication system is trained on an assumed stochastic channel model that is close to the actual channel effects at the first stage. Subsequently, the receiver is then finetuned based on the difference between the assumed stochastic channel model and the practical channel model at the second stage. This proposed scheme showcases the possibility of conducting over-the-air data transmission entirely carried out by NNs. However, the transmitter cannot be optimized during the finetuning of the receiver, which hinders the overall learning of the E2E communication system. There are some existing works studying to improve the two-step training approach, which can be classified into two main categories, i.e., channel imitation approaches and receiver-aided approaches.
@@ II-B Channel Imitation Schemes for End-to-End Communication Systems
In order to learn the channel effects in practical channel models and address the challenges of local optimization caused by the BP algorithm during the auto-encoder training [8], the channel imitation scheme using generative adversarial network (GAN) is widely studied [9]. In [10], a novel conditional GAN-based channel agnostic E2E communication system is developed to adaptively learn the channel effects of different types of channel models. More specifically, the generator is trained to approximate the conditional data distribution of the practical channel effects, and the discriminator is trained to distinguish the real signals from the generated fake signals. By conditioning the generator and the discriminator on the channel, the system can learn to adapt to a variety of channel conditions, even if the channel characteristics are unknown beforehand. This channel agnostic E2E system empowers the generator to serve as a bridge connecting the transmitter and the receiver. This enables the BP of error gradients from the receiver to the transmitter, ultimately facilitating the overall learning of the transceiver. This conditional GAN-based channel imitation scheme is further enhanced in [11], where the system can not only adapt to various channel models by learning their characteristics but can also transmit long sequences of information bits. In particular, this proposed system has a significantly extended block length, from 8 information bits to 128 information bits, without much degradation in BLER performance. However, the GAN-based schemes face challenges that they struggle to capture complex channel distributions and require extensive training data, leading to poor generalization in real-world channel conditions.
A residual-aided GAN (RA-GAN) based E2E communication system is proposed to solve the above challenges by applying residual neural network (ResNet) [12] to the generator to enable better generation performance for the received signals and adding $\bm{l}_{2}$ regularizer to punish large weights in E2E system [13]. However, this scheme experiments only on a block length of 4 bits, which falls considerably short of the practical scenarios for data transmission. Despite the effectiveness of the aforementioned channel imitation schemes in addressing the local optimization issue associated with the BP algorithm, GAN-based solutions still show limited performance with large block lengths. This limitation is particularly evident in more complex wireless channels, such as the Rayleigh fading channel. As such, further improvements are required to address these challenges and enhance the system’s performance.
@@ II-C Receiver-Aided Schemes for End-to-End Communication Systems
Recently, DRL has demonstrated its great potential in wireless communications and networking [14], particularly in the field of E2E communication systems design. Apart from the channel imitation solutions, RL techniques have been emerging as a promising tool in receiver-aided approaches, enabling data transmission without the need for explicit channel modelling or CSI [15]. In particular, the simultaneous perturbation stochastic approximation technique is used in [16] to train the transmitter, where the transmitter is optimized by utilizing the loss value obtained from the receiver. In [17], a combination of RL and supervised learning techniques is employed to optimize both the transmitter and the receiver. In [18], the authors extend these works by considering a noisy feedback loop. Particularly, the proposed system with a noisy feedback link demonstrates identical performance with the system with a perfect feedback link. In summary, the aforementioned receiver-aided solutions circumvent the unknown channel issue and local optimization problem arising from BP by using the loss value obtained from the receiver. However, the biggest issue of the receiver-aided solutions is that the variance of loss value from the receiver will scale with the increasing number of channel uses and the increasing block length of the input message, which leads to significant performance degradation and slow convergence. In addition, the training of communication systems over complex wireless communication channels faces challenges due to the relatively long coherence time where only small portions of CSI are learned, which leads to poor sample efficiency and also slow convergence. Given the above disadvantages of the channel imitation approaches and receiver-aided approaches, there is currently no effective solution that can transmit data with large block lengths (e.g., 256 information bits) over an unknown channel without compromising the BLER performance.
II-D CNN-based Architecture for Binary Encoding
CNNs are better suited than Multi-layer perceptrons (MLPs) for handling large block lengths with binary encoding because they can effectively capture local patterns through convolutional layers, which enhances their ability to recognize bit sequences. Additionally, CNNs benefit from parameter efficiency due to weight sharing, avoiding the parameter explosion issue that can occur with MLPs when dealing with large block sizes [19]. In order to tackle the encoding of source information with large block length, the binary encoding is adopted rather than the one-hot encoding in many existing works including [20], [11] and [21]. These works introduce novel CNN architectures to mitigate the dimensionality explosion associated with larger block sizes. Specifically, the first two works, [20] and [11], focus on using CNNs for binary encoding with large block lengths, which requires explicit signal processing including channel estimation and channel equalization. In addition, the authors in [22] introduce an innovative Turbo Autoencoder (TAE) architecture that combines the strength of Turbo coding and CNNs to effectively model the encoder and decoder processes in communication systems. In particular, the proposed TAE exhibits superior bit error rate performance under low SNR conditions when compared to the conventional coding schemes, including LDPC and polar codes. Moreover, it demonstrates significant performance improvement for large block sizes over the classical Convolutional-Autoencoder model presented in [5]. In contrast, [21] extends the use of CNNs to a fully differentiable end-to-end system that includes detection, synchronization, equalization, and decoding. In our work, we develop a CNN architecture inspired by [20] to tackle the challenges of large block sizes. Therefore, the synchronization and channel estimation are not considered in the work. With the proposed CNN, we effectively train an E2E communication system for longer block sizes without explicit information about channel distributions in advance.
@@ III System Model
In this section, we will introduce the basic concepts of E2E communication systems. As shown in Fig. 1(b), an auto-encoder-based E2E communication system consists of a transmitter, a receiver, and a mathematically formulated channel, where the transmitter and the receiver are represented as two independent NNs. The transmitter NN aims to encode the original binary message $\bm{m}$ to the encoded signal $\bm{x}\in{\mathbb{C}}^{n}$, where ${\mathbb{C}}^{n}$ denotes that the encoded signal $\bm{x}$ is presented as complex numbers, making $n$ discrete uses of channel. The encoded signal is then sent via the channel to the receiver. The received signal after channel distortion is denoted by $\bm{y}\in{\mathbb{C}}^{n}$. The receiver aims to decode the received signals and recover the original binary message $\bm{m}$. The implementations of the E2E communication system can be illustrated as the sequence of three individual functions, which can be expressed as:
@@
where $f_{\mathrm{E}}$ represents the encoder function mapping the original binary message to the encoded signal, i.e., $\bm{x}=f_{\mathrm{E}}(\bm{m};\bm{\theta}_{\mathrm{E}})$, and $\bm{\theta}_{\mathrm{E}}$ denotes the trainable parameters of the transmitter NN. $f_{\mathrm{h}}$ represents the channel impairments with channel realization $h$, i.e., $\bm{y}=f_{\mathrm{h}}(\bm{x})$. $f_{\mathrm{D}}$ represents the decoder function that aims to recover the received signals to the estimated message, i.e., $\hat{\bm{m}}=f_{\mathrm{D}}(\bm{y};\bm{\theta}_{\mathrm{D}})$, and $\bm{\theta}_{\mathrm{D}}$ denotes the trainable weights of the receiver NN. The goal of the E2E communication system is to jointly optimize both the transmitter NN and the receiver NN by minimizing the E2E loss function, denoted by $L=\mathcal{L}(\bm{m},\hat{\bm{m}})$. The loss function $\mathcal{L}(\bm{m},\hat{\bm{m}})$ is regarded as the objective function which measures the distance between the original message $\bm{m}$ and the estimated message $\hat{\bm{m}}$ for such E2E framework.
@@
The E2E learning approach considers the entire system as an auto-encoder and trains the NNs of the transceivers in a supervised manner. The training process involves calculating the error gradients of each layer by utilizing the BP algorithm. To enable the BP of the error gradients, it is necessary to ensure that the loss function with respect to the trainable weights at each layer of the NN is differentiable. The computed error gradients subsequently update the weights and bias at each layer, with the goal of minimizing the overall E2E loss. Therefore, the E2E learning approach entails a differentiable channel function beforehand to facilitate the BP of error gradients from the receiver to the transmitter. This enables the joint optimization of the entire E2E communication system. However, while the E2E learning paradigm demonstrates efficiency in simple channel models, its performance becomes notably constrained when confronted by more complex communication scenarios. The current limitations of the E2E learning paradigm can be summarized as follows. First, the primary limitation of the E2E learning method is its reliance on an explicit differentiable channel model, which is often impractical due to real-world communication challenges. Second, the non-differentiable channel model potentially hinders the gradient-based optimization, preventing the overall learning of the system. Third, obtaining real-time CSI is difficult in practice, undermining the validity of assumed channel functions. Furthermore, the existing literature primarily focuses on research with short block lengths, and there is a lack of evidence regarding the performance of solutions for block lengths exceeding 100 bits.
@@
This paper introduces a DDPG-based E2E communication system aimed at addressing the challenges posed by non-differentiable channel models and imperfect CSI while improving the performance in terms of data transmission with long block lengths. By leveraging the DDPG algorithm, the loss value of the receiver NN serves as feedback information for updating the transmitter NN. Hence, the transmitter NN and the receiver NN can be implicitly and jointly optimized without requiring prior knowledge of channel models. The following sections will present a comprehensive analysis of our proposed DDPG-based E2E solution.